# Phase transition like behaviors of Propagation of Passenger Stranding phenomena in Subway Networks

### 12 Jan 2024 v1

@@ 2.1 System I: Steady-State features
Assuming spatial and temporal homogeneity in both supply and demand, the system can reach a non-equilibrium steady state. By studying systems in this steady state, the influence of the subway network’s topological structure on the PPS phenomenon can be revealed. In System I, supply scale remains constant by establishing a consistent train departure interval (4 minutes) from the originating station. The supply-demand ratio in the system by adjusting the demand scale, $P_{u}$ (the number of passengers entering the station per minute). The specific simulation methodology can be found in the Supplement Meterials, A.2.3. After a relaxation period of finite duration, the system reaches a non-equilibrium steady state, where the average waiting time of passengers and the values of these 3 order parameters steadied at specific values (examples in Supplement Meterials, B.2.1, Fig.B.13). The time dependent number of stranded passengers can be seen in Supplement Meterials, B.2.2, Fig.B.14, Fig.B.15. System responsive behavior to the demand scale, $P_{u}$, can be characterized by the time-averaged waiting time $t_{w}$ and the 3 order parameters that describe the scale of stranding, which is shown in Fig.2. Both $t_{w}$ and 3 order parameters exhibit similar responsive behaviors with a transition point around $P_{u}$=20. When $P_{u}$ is small, there are few stations with stranded passengers on the platforms randomly, and passengers can board the trains after a short waiting time. There is no spatial-temporal correlation between stations with stranded passengers, indicating the absence of the PPS phenomenon. Therefore, all 3 order parameters that characterize stranding scale tend to zero at small $P_{u}$ values. When $P_{u}$>20, $t_{w}$ increases abruptly with $P_{u}$, indicating a significant decrease in system efficiency. Stations with stranded passengers start to correlate with each other, forming clusters. The stranding scales indicated by all these 3 order parameters rapidly increase with increasing $P_{u}$. The behavior that significant changes in its response properties beyond a certain threshold is similar to that of a continuous phase transition. The threshold value, $P_{u}$=20, can be regarded as the transition point of this phase transition. When $P_{u}$ is below this threshold, there is no correlation in the occurrence of stranding phenomena between the stations. We call this state as a normal phase. However, when $P_{u}$ exceeds this threshold, stranded stations correlate with each other, forming persistent clusters of stations with stranded passengers that are difficult to dissipate. We call this state as a stranded phase. The system exhibits an abrupt change in its response behaviors to the demand scale exceeds the transition point, and the service efficiency dramatic decreased. We can use the value of $P_{u}$ at transition point as the threshold beyond which the system can no longer provide normal service.
@@
Figure 2: (A) Average waiting time of passengers, $t_{w}$, as a responsive behavior of the demand scale, $P_{u}$. (B)-(D) show the curves illustrating the changes in n_stranded, max_cluster, and tot_congestion as responsive behaviors of $P_{u}$. A phase transition like behavior occurs at $P_{u}=20$.
@@
Because there is no spatiotemporal matching issue between demand and supply in steady-state conditions, the formation of stranding and taking place of PPS phenomena solely arise from topological factors. The topological factor is characterized by the network betweenness centrality of each station. We find that when system is in the stranded phase, most of the stranded stations tend to cluster around the stations with high network betweenness centrality. To depict the dependence between the occurrence of stranding at stations on the topological factor, we sort the stations in stranded phase (taking $P_{u}$=25, 30, 35, and 40 as examples) in descending order based on the number of stranded passengers and the betweenness centrality, respectively. Using the ranks of these 2 queues as the vertical and horizontal coordinates in Fig. 3. If all stations fall on the diagonal line of this graph, the stranding phenomena and topological factors are linearly correlated. As indicated by Fig. 3 when systems are in stranded phase, the stations with higher betweenness centrality exhibit relatively higher numbers of stranded passengers and those stations with lower betweenness centrality have no passenger stranded. Taking $P_{u}$=30 as an example (blue circle), there are 33 stranded stations, with the maximum value of the vertical axis representing the ranking of passenger strand being 34. Those stations with rank above 74 which have lower betweenness centrality do not experience passenger stranding.
@@
Generally, a station with higher betweenness centrality indicates a higher level of connectivity with the network, which would reduce the occurrence of stranding. However, results in Fig. 3 show contrary to intuition. Indeed, this outcome reflects the characteristics of subway service. In systems under steady-state condition, since the passenger demands (pairs OD stations) are randomly generated in space, it has a high probability to require transfers to different lines at stations with higher betweenness centrality. Although these stations have high connectivity, passengers have to wait for the transfer train to arrive before they can continue their travel, resulting in a delay in receiving service. It should be noted that train scheduling in different lines is independent of each other, so passengers needing transfers generally cannot receive immediate service. This is equivalent to increasing the demand volume at the transfer stations acting as hubs, leading to stranding. By analyzing the steady-state conditions, the emergence stranding phenomena and formation of clusters in the network are observed, which shows a phase transition like behavior. The mechanisms behind this behavior are the network’s topology and the independence of scheduling in different lines in subway. The steady-state conditions analyzing is used to reveal the feature of spontaneous emergence of stranded cluster in subway. However, such phenomena are not severe in real subway service, as the OD pairs of actual commuting passengers are not random. Citizens choose residence and workplace to minimize the possibility of transfers and enhance travel convenience. Hereafter, to investigate the PPS phenomenon in real systems, we will establish a computational model based on Automatic Fare Collection (AFC) data.
@@
Figure 3: Relationship between the ranking of station betweenness centrality and the ranking of stranded passenger numbers. The betweenness centrality ranking of stations is based on the descending order of the betweenness centrality values for each station.
2.2 System II: Spatio-temporal Imbalance between Supply and Demand
@@
Large cities exhibit a large spatial separation between job and residence centers, particularly influenced by the TOD urban development model. Citizens tend to select residence and job locations along subway lines, resulting in a spatiotemporal imbalance in commuting demands with rush hours and long-distance travel patterns. To model the non-steady-state evolution of PPS, a tidal supply-demand model at a daily time scale based on the spatiotemporal characteristics of AFC data and train timetable data is constructed. The AFC data covers the entire weekdays from May 10th to May 14th, 2021, for the nine subway lines depicted in Fig.1B of the model (more details in Supplement Meterials, A.2). The AFC data consists of 3,811,418 records of passengers ($N_{r}$) and 5,293 records of trains ($N_{ori}$). The timetable data is from Beijing Rail Transit App (more details in Supplement Meterials, A.1). The spatiotemporal mismatch between supply and demand is the primary mechanism of PPS phenomena. In the simulation, the supply scale is fixed at the real supply level while varying the demand scale to investigate the effects of service shortages on PPS. The demand scale is defined as $P_{r}=N_{d}/N_{r}$, where Nd represents the total number of passengers taking subway throughout the day in the simulation, and $N_{r}$ represents the real total number of passengers. We generate Nd OD pairs of demand by sampling the OD pairs in the same time interval based on the probability $P_{r}=N_{d}/N_{r}$ from the real data throughout the weekdays (see Supplement Meterials, A.2.4 and Fig. A.12 for parameter settings and validation).
@@
Unlike the results obtained under steady-state conditions, the computational results based on real-system data exhibit spatiotemporal heterogeneity. In spatial dimension, there are 4 significant spatially independent clusters of stranded stations (Cluster I-IV), regardless of the demand scale, whether it is low ($P_{r}=1.0$) (Supplement Meterials, B.3.3, Fig.B.17(E) or high ($P_{r}=1.6$) (Supplement Meterials, B.3.3, Fig.B.17(J)). These clusters represent distinct commuting "corridors" connecting the residence centers located in the suburb (north: Huilongguan, Tiantongyuan; south: Daxing; east: Beijing Municipal Administrative Center and North Three Counties, including: Sanhe City, Dachang Hui Autonomous County and Xianghe County; west: Shijingshan) to the job centers situated in the city center. The formation of these commuting corridors is attributed to the mutual influence of subway construction on the spatial distribution of residential and employment areas, thereby characterizing Beijing as a Transit-Oriented Development (TOD) type city. The list contains the stations (names and labels) included in each Cluster can be found in Supplement Meterials, B.3.3, Tab.B.6,B.7. In temporal dimension, instead of steady-state features, the system exhibits characteristics of rush hours. Several time points are selected to observe PPS, as depicted in Supplement Meterials, B.3.3, Fig.8. For a demand scale of $P_{r}=1.0$, at 7:10, only a few isolated stations manifest stranding phenomena, and these stations are spatially independent. By 8:00, PPS occurs, signifying the association of stations with stranded passengers and the formation of clusters. By 8:20, the size of these clusters reach their maximum. The stranding phenomena completely vanish from the network by 13:00 (Supplement Meterials, B.3.3, Fig.8(A)-(D)). When the demand scale is high, such as $P_{r}=1.6$ (Supplement Meterials, B.3.3, Fig.8(F)-(I)), with the arrival of the morning rush hour, the system demonstrates associations between stations with stranded passengers, leading to cluster formation. As the morning rush hour subsides, the scale of PPS gradually decreased. In contrast to the case with a lower demand scale, during the onset of the evening rush hour at 13:00, the clusters generated during the morning rush hour do not entirely dissipate (Supplement Meterials, B.3.3, Fig.8(I)). The system is always in the working state with high stranding phenomenon, and the efficiency of service delivery is quite low.
@@
In the context of spatiotemporal heterogeneity driven by real-system data, both passenger waiting time and 3 order parameters fail to reach steady states, exhibiting dynamic fluctuations over time. Consequently, the method to determining the transition point of the system, as depicted in Figure 2, becomes unfeasible. Alternatively, to characterize the non-equilibrium responsive behaviors of the system, one can introduce non-steady-state dynamics by increasing the demand level in a system that is in a steady state and measuring the relaxation time required for the system to establish a new steady state. In this study, the morning rush hour and evening rush hour (Supplement Meterials, B.3.1, Fig.B.16) are considered as two successive stimuli imposed on the system. Prior to the morning rush hour, the system is in a steady state where no stations with stranded passengers are observed. If the relaxation time of the system after the stimulus of morning rush hour is shorter than the time interval between the morning rush hour and evening rush hour, the system consistently returns to a steady state before the evening rush hour begins. This state is referred to as the normal state in present work. On the other hand, if the relaxation time is shorter than the time interval between the morning rush hour and evening rush hour, the accumulation of stranding phenomena within the system leads to a continuous decline in service efficiency. This state is referred to as the stranded state. By varying the passenger demand scale ($P_{r}$), the changes in average passenger waiting time and the 3 order parameters over time are plotted to observe the system’s responsive behaviors, as presented in Figure 4. The average passenger waiting time directly reflects the system’s service efficiency. To illustrate the phase transition characteristics, Figure 4A presents the summation of the average passenger waiting time over the entire day as a function of $P_{r}$. Importantly, a similar transition point, as shown in Figure 2A, is observed at $P_{r}\approx 1.4$, which can be defined as the threshold for resilience failure. When the demand scale is below this threshold, such as when $P_{r}=1.0$ representing the demand scale in a real normal day, 3 order parameters reflecting the level of stranded passengers exhibit distinct morning and evening rush hour characteristics with 2 peaks in the temporal dimension. Moreover, these 2 peaks do not overlap, creating a clear time window where no stranding phenomenon occurs. This finding implies that under normal demand scale for the morning rush hour and evening rush hour stimulus, the existing subway service effectively suppresses the occurrence of PPS within a limited time, without correlating to the subsequent rush hour. As the demand scale decreases, both the peak values of the 3 order parameters for the morning rush and evening hour diminish. When $P_{r}<1.0$, only the morning rush hour exhibits stations with stranded passengers and stranding clusters, while the evening rush hour is nearly absent. As $P_{r}$ increases, the relaxation time for the morning rush hour extends, and the start time of the evening rush hour is advanced. After $P_{r}$ reaches 1.4, the stranded station clusters generated by the morning rush hour fail to dissipate completely before the onset of the evening rush hour. Consequently, there is no time window without stranding between the morning rush hour and evening rush hour, making the subway network incapable of providing effective service due to PPS. This observation contrasts with the distinct dynamic characteristics of the morning rush and evening hour that are clearly separated when $P_{r}<1.4$. Hence, we define $P_{r}=1.4$ as the transition point of this system, serving as a metric for measuring threshold of service resilience failure.
Figure 4: (A) Average passenger waiting time ($t_{w}$) as a function of $P_{r}$. (B)-(D) Temporal evolution of max_cluster, n_stranded, and tot_stranded throughout the day as functions of $P_{r}$.
2.2.1 Dynamics Mechanism of Cluster Growth

### 28 Feb 2024 v2

@@ 2.1 System I: Steady-state features
Assuming spatial and temporal homogeneity in both supply and demand, the system can reach a non-equilibrium steady state. By studying systems in this steady state, the influence of the subway network’s topological structure on the PPS phenomenon can be revealed. In System I, supply scale remains constant by establishing a consistent train departure interval (4 min) from the originating station. The supply-demand ratio in the system by adjusting the demand scale, $P_{\mathrm{u}}$ (the number of passengers entering the station per minute). The specific simulation methodology can be found in the Supplementary Information, A.2.3. After a relaxation period of finite duration, the system reaches a non-equilibrium steady state, where the average waiting time of passengers and the values of these 3 order parameters steadied at specific values (examples in Supplementary Information, B.2.1, Fig. B.4). The time dependent number of stranded passengers can be seen in Supplementary Information, B.2.2, Fig. B.5, Fig. B.6. System responsive behavior to the demand scale, $P_{\mathrm{u}}$, can be characterized by the waiting time $t_{\mathrm{w}}$, $n_{\mathrm{s}}$, $m_{\mathrm{c}}$ and $n_{\mathrm{t}}$ that are averaged with respect to time within steady state, which is shown in Fig. 2. Both $t_{\mathrm{w}}$ and the 3 order parameters exhibit similar responsive behaviors with a transition point around $P_{\mathrm{u}}=20$ where the slopes of these 4 lines in Fig. 2 show observable changes. When $P_{\mathrm{u}}$ is small, there are few stations with stranded passengers on the platforms randomly, and passengers can board the trains after a short waiting time. There is no spatial-temporal correlation between stations with stranded passengers, indicating the absence of the PPS phenomenon. Therefore, all 3 order parameters that characterize stranding scale tend to zero at small $P_{\mathrm{u}}$ values. When $P_{\mathrm{u}}>$20, $t_{\mathrm{w}}$ increases abruptly with $P_{\mathrm{u}}$, indicating a significant decrease in system efficiency. Stations with stranded passengers start to correlate with each other, forming clusters. The stranding scales indicated by all these 3 order parameters rapidly increase with increasing $P_{\mathrm{u}}$. The behavior that significant changes in its response properties beyond a certain threshold is similar to that of a continuous phase transition. The threshold value, $P_{\mathrm{u}}$ = 20, can be regarded as the transition point of this phase transition like behavior. When $P_{\mathrm{u}}$ is below this threshold, there is no correlation in the occurrence of stranding phenomena between the stations. We refer to this state as the normal phase. When $P_{\mathrm{u}}$ exceeds this threshold, stranded stations correlate with each other, forming persistent clusters of stations with stranded passengers that are difficult to dissipate. We refer to this state as the stranded phase. In this phase, the system can no longer provide normal service. Therefore, We can use the value of $P_{\mathrm{u}}$ at transition point as the threshold of resilience failure.
@@
Figure 2: a Average waiting time of passengers, $t_{\mathrm{w}}$, as a responsive behavior of the demand scale, $P_{\mathrm{u}}$. b-d show the curves illustrating the changes in $n_{\mathrm{s}}$, $m_{\mathrm{c}}$, and $n_{\mathrm{t}}$ as responsive behaviors of $P_{\mathrm{u}}$. A phase transition like behavior occurs at $P_{\mathrm{u}}=20$.
@@
Because there is no spatiotemporal matching issue between demand and supply in steady-state conditions, the formation of stranding and taking place of PPS phenomena solely arise from the topological factor. The topological factor is characterized by the network betweenness centrality of each station. We find that when system is in the stranded phase, most of the stranded stations tend to cluster around the stations with high network betweenness centrality. To depict the dependence between the occurrence of stranding at stations on the topological factor, we sort the stations in stranded phase (taking $P_{\mathrm{u}}$ = 25, 30, 35, and 40 as examples) in descending order based on the number of stranded passengers and the betweenness centrality, respectively. Using the ranks of these 2 queues as the vertical and horizontal coordinates in Fig. 3. If all stations fall on the diagonal line of this graph, the stranding phenomena and topological factors are linearly correlated. As indicated by Fig. 3, when systems are in the stranded phase, the stations with higher betweenness centrality exhibit relatively higher numbers of stranded passengers and those stations with lower betweenness centrality have no passenger stranded. Taking $P_{\mathrm{u}}$ = 30 as an example (blue circle), there are 33 stranded stations, with the maximum value of the vertical axis representing the ranking of stranded passengers being 34. Those stations with rank above 74 which have lower betweenness centrality do not experience passenger stranding.
@@
      Generally, a station with higher betweenness centrality indicates a higher level of connectivity with the network, which would reduce the occurrence of stranding. However, results in Fig. 3 show contrary to intuition. Indeed, this outcome reflects the characteristics of subway service. In systems under steady-state condition, since the passenger demands (pairs OD stations) are randomly generated in space, it has a high probability to require transfers to different lines at stations with higher betweenness centrality. Although these stations have high connectivity, passengers have to wait for the transfer train to arrive before they can continue their travel, resulting in a delay in receiving service. It should be noted that train scheduling in different lines is independent of each other, so passengers needing transfers generally cannot receive immediate service. This is equivalent to increasing the demand volume at the transfer stations acting as hubs, leading to stranding. By analyzing the steady-state conditions, the emergence of stranding phenomena and formation of clusters in the network are observed, which shows a phase transition-like behavior. The mechanisms underlying this behavior include the subway network’s topology and the independent scheduling of different subway lines. The steady-state conditions analyzing is used to reveal the feature of spontaneous emergence of stranded clusters in the subway. However, such phenomena are not severe in real subway service, as the OD pairs of actual commuting passengers are not random. Citizens choose residence and workplace to minimize the possibility of transfers and enhance travel convenience. Hereafter, to investigate the PPS phenomenon in real systems, we will establish a computational model based on Automatic Fare Collection (AFC) data which provides the real OD pairs depending on time.
@@
Figure 3: Relationship between the ranking of station betweenness centrality and the ranking of stranded passenger numbers which shows a positive correlation. Both 2 rankings are obtained through sorting in descending order based on the betweenness centrality values and the passenger number for each station.
2.2 System II: Spatiotemporal imbalance between supply and demand
@@
Large cities exhibit a large spatial separation between job and residence centers, particularly influenced by the TOD urban development model. Citizens tend to select residence and job locations along subway lines, resulting in a spatiotemporal imbalance in commuting demands with rush hours and long-distance travel patterns. To model the non-steady-state evolution of PPS, a tidal supply-demand model at a daily time scale based on the spatiotemporal characteristics of AFC data and train timetable data is constructed. The AFC data covers the entire weekdays from May 10th to May 14th, 2021, for the 9 subway lines depicted in Fig. 1b of the model (more details of passengers’ related data in Supplementary Information, A.1.1). The AFC data consists of 3,860,008 records of passengers ($N_{\mathrm{r}}$) and 5,266 records of trains ($N_{\mathrm{ori}}$). The timetable data is from Beijing Rail Transit App (more details in Supplementary Information, A.1.2). The spatiotemporal mismatch between supply and demand is the primary mechanism of PPS phenomena. In the simulation, the supply scale is fixed at the real supply level while varying the demand scale to investigate the effects of service shortages on PPS. The demand scale is defined as $P_{\mathrm{r}}=N_{\mathrm{d}}/N_{\mathrm{r}}$, where $N_{\mathrm{d}}$ represents the total number of passengers taking subway throughout the day in the simulation, and $N_{\mathrm{r}}$ represents the real total number of passengers. We generate $N_{\mathrm{d}}$ OD pairs of demand by sampling the OD pairs in the same time interval based on the probability $P_{\mathrm{r}}=N_{\mathrm{d}}/N_{\mathrm{r}}$ from the real data throughout the weekdays (see Supplementary Information, A.2.4 and Fig. A.3 for parameter settings and validation).
@@
Unlike the results obtained under steady-state conditions, the computational results based on real-system data exhibit spatiotemporal heterogeneity. The time dependent number of stranded passengers can be observed in Supplementary Movie 1 for $P_{\mathrm{r}}=1.0$, and Supplementary Movie 2 for $P_{\mathrm{r}}=1.6$. In spatial dimension, there are 4 significant spatially independent clusters of stranded stations (Cluster I-IV), regardless of the demand scale, whether it is low ($P_{\mathrm{r}}=1.0$) (Supplementary Information, B.3.3, Fig. B.8e) or high ($P_{\mathrm{r}}=1.6$) (Supplementary Information, B.3.3, Fig. B.8j). These clusters represent distinct commuting “corridors” connecting the residence centers located in the suburb (north: Huilongguan, Tiantongyuan; south: Daxing; east: Beijing Municipal Administrative Center and North Three Counties, including: Sanhe City, Dachang Hui Autonomous County and Xianghe County; west: Shijingshan) to the job centers situated in the city center. The formation of these commuting corridors is attributed to the mutual influence of subway construction on the spatial distribution of residential and job areas, thereby characterizing Beijing as a TOD type city. The list containing the stations (names and labels) included in each cluster can be found in Supplementary Information, B.3.3, Table B.6 and B.7. In temporal dimension, instead of steady-state features, the system exhibits characteristics of rush hours. Several time points are selected to observe PPS, as depicted in Supplementary Information, B.3.3, Fig. B.8. For a demand scale of $P_{\mathrm{r}}=1.0$, at 7:10, only a few isolated stations manifest stranding phenomena, and these stations are spatially independent. By 8:00, PPS occurs, signifying the association of stations with stranded passengers and the formation of clusters. By 8:20, the size of these clusters reaches their maximum. The stranding phenomena completely vanish from the network by 13:00 (Supplementary Information, B.3.3, Fig. B.8a-d). When the demand scale is high, such as $P_{\mathrm{r}}=1.6$ (Supplementary Information, B.3.3, Fig. B.8f-i), with the arrival of the morning rush hour, the system demonstrates associations between stations with stranded passengers, leading to cluster formation. As the morning rush hour subsides, the scale of PPS gradually decreased. In contrast to the case with a lower demand scale, during the onset of the evening rush hour at 13:00, the clusters generated during the morning rush hour do not entirely dissipate (Supplementary Information, B.3.3, Fig. B.8i). The system is always in the working state with high stranding phenomenon, and the efficiency of service delivery is quite low.
@@
In the context of spatiotemporal heterogeneity driven by real-system data, both passenger waiting time and 3 order parameters fail to reach steady states, exhibiting dynamic fluctuations over time. Consequently, the method to determining the transition point of the system, as depicted in Fig. 2, becomes unfeasible. Alternatively, to characterize the non-equilibrium responsive behaviors of the system, one can introduce non-steady-state dynamics by increasing the demand level in a system that is in a steady state and measuring the relaxation time required for the system to establish a new steady state. In this study, the morning rush hour and evening rush hour (Supplementary Information, B.3.1, Fig. B.7) are considered as 2 successive stimuli imposed on the system. Prior to the morning rush hour, the system is in a steady state where no stations with stranded passengers are observed. If the relaxation time of the system after the stimulus of morning rush hour is shorter than the time interval between the morning rush hour and evening rush hour, the system consistently returns to a steady state before the evening rush hour begins. This state is referred to as the “normal phase” in present work. Conversely, if the relaxation time is shorter than the time interval between the morning rush hour and evening rush hour, the accumulation of stranding phenomena within the system leads to a continuous decline in service efficiency. This state is referred to as the stranded phase. By varying the passenger demand scale ($P_{\mathrm{r}}$), the changes in average passenger waiting time and the 3 order parameters over time are plotted to observe the system’s responsive behaviors, as presented in Fig. 4. The average passenger waiting time directly reflects the system’s service efficiency. To illustrate the phase transition characteristics, Fig. 4a presents the summation of the average passenger waiting time over the entire day as a function of $P_{\mathrm{r}}$. Importantly, a similar transition point, as shown in Fig. 2a, is observed at $P_{\mathrm{r}}\approx 1.4$, which can be defined as the threshold for resilience failure. When the demand scale is below this threshold, such as when $P_{\mathrm{r}}=1.0$ representing the demand scale in a real normal day, 3 order parameters reflecting the level of stranded passengers exhibit distinct morning and evening rush hour characteristics with 2 peaks in the temporal dimension. Moreover, these 2 peaks do not overlap, creating a clear time window where no stranding phenomenon occurs. This finding implies that under normal demand scale for the morning rush hour and evening rush hour stimulus, the existing subway service effectively suppresses the occurrence of PPS within a limited time, without correlating to the subsequent rush hour. As the demand scale decreases, both the peak values of the 3 order parameters for the morning rush and evening hour diminish. When $P_{\mathrm{r}}<1.0$, only the morning rush hour exhibits stations with stranded passengers and stranding clusters, while the evening rush hour is nearly absent. As $P_{\mathrm{r}}$ increases, the relaxation time for the morning rush hour extends, and the start time of the evening rush hour is advanced. After $P_{\mathrm{r}}$ reaches 1.4, the stranded station clusters generated by the morning rush hour fail to dissipate completely before the onset of the evening rush hour. Consequently, there is no time window without stranding between the morning rush hour and evening rush hour, making the subway network incapable of providing effective service due to PPS. This observation contrasts with the distinct dynamic characteristics of the morning rush and evening hour that are clearly separated when $P_{\mathrm{r}}<1.4$. Hence, we define $P_{\mathrm{r}}=1.4$ as the transition point of this system, serving as a metric for measuring the threshold of service resilience failure.
Figure 4: a Averaged passenger waiting time ($t_{\mathrm{w}}$) as a function of $P_{\mathrm{r}}$ which gives a transition point around $p_{\mathrm{r}}=1.4$. b-d Temporal evolution of $m_{\mathrm{c}}$, $n_{\mathrm{s}}$, and $n_{\mathrm{t}}$ throughout the day as functions of $P_{\mathrm{r}}$ which show rush hours of morning and evening. The transition from Normal phase to Stranded phase can be determined by whether there is a correlation between the 2 peak periods.
2.3 Dynamics mechanism of cluster growth